RE: revisions you made to TOF Systems last access

Date view Thread view Subject view Author view Attachment view

From: Xu, Zhangbu (xzb_at_bnl.gov)
Date: Mon Feb 24 2003 - 09:55:38 EST


 

-----Original Message-----
From: MAJKA_at_hepmail.physics.yale.edu
Cc: Xu, Zhangbu
Sent: 2/24/03 8:03 AM
Subject: Re: revisions you made to TOF Systems last access

Hi Bill;
  Although I was directly involved only at the back
end (ie where the signals come to the trigger racks),
my understanding is that only the "taps" to 69/72
signals from TOFr were plugged in. Zhangbu can probably
add more details. It is also true that Jeff Landgraff
did a lot of work on the DAQ end "recompiled all config.
file" or something like that. At least one thing (TPC
laser trigger) seemd to need some fixing after that.
I'll be at BNL later today, and for the rest of the
week, so if there is anything I or Zhangbu can do to
help find the problem, please ask.
  One other tidbit of information, we did plug the
69-fold OR from TOFr into a chan. of the RICH scalers,
and at least on Friday (beam off) it was counting at
about 200Hz - which sound like a reasonable noise rate.
Our local WEB server at Yale is down, so I can't get to
the RICH scaler db, but later today when I am at BNL
I can take a look and see if at least this scaler sees
the beam. Again, please let me know if I can be of any
help in finding the problem.
     best regards, Dick Majka.


Date view Thread view Subject view Author view Attachment view

This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:35 EDT